# Talha Bozkus and Urbashi Mitra

### v2

@@ III-2 State
The individual state of TXi is a triplet $s_{i}=(x_{i},y_{i},I_{i})$, where $(x_{i},y_{i})$ is its location with $x_{i},y_{i}\in[0,\Delta_{L},2\Delta_{L},...,L]$, and $I_{i}$ is the ARSS. Since ARSS may be continuous, we quantize $I_{i}$ to the nearest value in $[I_{\text{min}},I_{\text{max}}]$, which is discretized into equal steps of size $\Delta_{I}$. The transition from $(x_{i},y_{i})$ to $(x_{i}^{\prime},y_{i}^{\prime})$ depends on the random movement of TXi, while the transition from $I_{i}$ to $I_{i}^{\prime}$ is influenced by several factors such as the random movement of all TXs, the distances between TXs, $L$, and $I_{thr}$. The joint state $\bar{s}$ is the concatenation of individual states as $\bar{s}=(s_{1},s_{2},...,s_{N_{T}})$, and is not available to the agents (i.e. agents do not have access to each others’ states).
The state-space is divided into two parts: coordinated and uncoordinated states. The joint state $\bar{s}$ is coordinated if there is at least one strong ARSS measurement in the wireless network, i.e., there is at least one agent $i$ such that $I_{i}>I_{thr}$, where $I_{thr}$ is the predefined ARSS threshold; otherwise, it is uncoordinated. Thus, if at least one TX identifies the state as coordinated, it informs TX1, which then broadcasts this information to all TXs. We denote $|\mathcal{S}_{C}|$ and $|\mathcal{S}_{U}|$ as the number of coordinated and uncoordinated states, which depend on $I_{thr}$, and $|\mathcal{S_{C}}|+|\mathcal{S_{U}}|=|\mathcal{S}|$.
We note that there are alternative approaches to determining coordinated states, such as based on predefined states [13], proximity between agents [40], and statistical analysis of prior trajectories [41]. Our approach is similar to [42], which also uses local RSS measurements to determine coordinated states. However, in our work, coordination is based on sufficiently high ARSS values. In addition, coordination is possible even if not all agents measure high ARSS levels. This makes our scheme robust against ARSS measurement inaccuracies in some agents, and scalable by reducing dependency on each agent’s measurement.
@@ III-3 Action
The action set of TXi is $\mathcal{A}_{i}=\{1,2,\ldots,N_{B}\}$, where each index corresponds to a specific BS that the TX may connect. If a TX is outside a BS’s coverage area, the action is invalid and assigned an infinite cost to prevent selection. Thus, $|\mathcal{A}_{i}|\leq N_{B}$. The joint action $\bar{a}$ is the concatenation of individual actions as $\bar{a}=(a_{1},a_{2},...,a_{N_{T}})$, and $|\mathcal{A}|\leq{N_{B}}^{N_{T}}$.
@@ III-4 Cost function
$c_{i}=\beta_{1}\frac{P_{i}}{\log_{2}(1+\operatorname{SNR}_{i})}+\beta_{2}Q(\sqrt{\text{SNR}_{i}})+\beta_{3}Q(\frac{I_{thr}-I_{i}}{\sigma}),$ (5)
where $\operatorname{SNR}_{i}=\frac{P_{i}}{d_{i}^{2}(n+I_{i})}$, $d_{i}$ is distance between TXi and the BS it is connected to, $P_{i}$ is the transmission power of TXi, and $n$ is the zero-mean Gaussian noise as $n\sim\mathcal{N}(0,\sigma^{2})$. Herein, the first component is the efficiency cost. A higher SNR with constant transmitting power improves the effective data rate and reduces the efficiency cost. The second component is the reliability cost. The $Q$-function represents the inverse relationship between SNR and bit error rate. Thus, higher SNR reduces the bit error rate and improves transmission reliability. The third component is the fairness cost. If the measured local ARSS exceeds the threshold $I_{thr}$, it suggests that some TXs are close to each other and may compete for the same resources (time slots, frequencies), potentially overloading some BSs while underutilizing others, which increases the fairness cost. The constants $\beta_{1},\beta_{2},\beta_{3}$ are the cost weights such that $\beta_{1}+\beta_{2}+\beta_{3}=1$. The joint cost is the sum of individual costs: $\bar{c}=\sum_{i=1}^{N_{T}}c_{i}(\bar{s}_{t},\bar{a}_{t})$. We emphasize that the individual cost function of each TX in (5) relies solely on local ARSS observations and does not depend on the knowledge of other TXs.
This cost function has several key properties. It is non-negative and bounded. It is highly non-linear and non-monotonic in some parameters (e.g., $P_{i}$), but monotonic in others (e.g., SNR and $I_{thr}$). Its non-convexity (due to the Gaussian $Q$-function) presents optimization challenges. The effect of different costs and parameters can be controlled through the weights $\beta_{i}$.
@@ IV-A Preliminaries and initialization
Inputs: $l$, $\gamma$, $\alpha_{t}$, $\epsilon_{t}$, $u_{t}$, $\mathbf{b}_{i,0}$, $Q_{i}^{(n)}$, $Q^{e}_{i}$, wireless parameters
@@
6:         Run MEMQ with real and synthetic environments
7:         Determine if current state is coordinated based on local ARSS, and if so, estimate it using the Bayesian approach
8:         Update the $Q$-functions as in Table II
@@
10:end for
11:$\bar{\pi}(\bar{s})\leftarrow\operatorname*{argmin}_{\bar{a}^{\prime}}\bar{Q}(\bar{s},\bar{a}^{\prime})$
The inputs to Algorithm 1 are algorithm parameters (trajectory length $l$, discount factor $\gamma$, learning rate $\alpha_{t}$, exploration probability $\epsilon_{t}$, update ratio $u_{t}$, initial belief vectors $\mathbf{b}_{i,0}$ for agent $i$ at $t=0$, number of iterations $T$) and wireless network parameters ($L$, $\Delta_{L}$, $N_{T}$, $N_{B}$, $r_{j}$, $I_{min}$, $I_{max}$, $\Delta_{I}$, $\beta_{1},\beta_{2},\beta_{3}$, $P_{i}$, $k$, $I_{thr}$, $\sigma$). The outputs are the joint $Q$-table, $\bar{Q}$, and the corresponding joint policy, $\bar{\pi}$.
@@ IV-B Creating synthetic environments
Each TXi independently interacts with the real environment Env1, collects samples $\{s_{i},a_{i},s_{i}^{\prime},c_{i}\}$, and stores this data in their buffer $\mathcal{B}_{i}$. Then, each TXi randomly samples from $\mathcal{B}_{i}$ to update their individual estimated PTT for Env1 ($\mathbf{\hat{P}}_{i}$), and simultaneously builds and updates the PTT of its synthetic environment ($\mathbf{\hat{P}}_{i}^{n}$) by taking the matrix power of $\mathbf{\hat{P}}_{i}$ as explained in Sec.II-C. We emphasize that creating synthetic environments is not a one-time process; the PTTs of synthetic environments are continuously updated as each TX continues sampling from the real environment. Thus, the accuracy of the estimated synthetic environments improves over time.
@@
IV-C Running MEMQ algorithms
While each TXi continuously updates its estimates for the PTT of real and synthetic environments, it also simultaneously runs a single-agent MEMQ algorithm using its two environments and continuously updates its local ensemble $Q$-functions $Q_{i}^{e}$. In other words, MEMQ algorithms are run on environments whose PTTs are continuously updated. This section of the algorithm is given in line 6.
The inputs to Algorithm 1 serve as inputs for the MEMQ algorithm of each TX. For example, $P_{i}$ varies across different MEMQs, but $I_{min}$ remains constant. While most inputs remain constant across different MEMQs, the different underlying PTTs and inherent randomness in sampling and exploration may lead to different convergence rates. As a result, this step of the algorithm is dominated by the slowest MEMQ.
We propose different single-agent MEMQ algorithms in our prior work [9, 8, 7, 5, 6]. While these algorithms share similar objectives and algorithm design, they have different assumptions and constraints on the underlying MDP and network parameters. We adopt the algorithm in [5] as it has the broadest applicability with minimal constraints and assumptions. For different wireless networks with specific assumptions or constraints, one may also employ the algorithms of [9, 8, 7, 6].
@@ IV-C Running MEMQ algorithms
U $\rightarrow$ C $Q_{i}^{(n)}(s_{i},a_{i})\leftarrow(1-\alpha)Q_{i}^{(n)}(s_{i},a_{i})+\alpha\big{[}c_{i}(s_{i},a_{i})+\frac{\gamma}{3}\max_{\bar{a}^{\prime}}\bar{Q}(\bar{s}^{\prime},\bar{a}^{\prime})\big{]}$,  $Q^{e}_{i}(s_{i},a_{i})\leftarrow$ MEMQ
C $\rightarrow$ U $\bar{Q}(\bar{s},\bar{a})\leftarrow(1-\alpha)\bar{Q}(\bar{s},\bar{a})+\alpha\sum_{i=1}^{3}\big{[}c_{i}(\bar{s},\bar{a})+\gamma\max_{a_{i}^{\prime}}Q_{i}^{e}(s_{i}^{\prime},a_{i}^{\prime})\big{]}$
C $\rightarrow$ C $\bar{Q}(\bar{s},\bar{a})\leftarrow(1-\alpha)\bar{Q}(\bar{s},\bar{a})+\alpha\big{[}\sum_{i=1}^{3}c_{i}(\bar{s},\bar{a})+\gamma\max_{\bar{a}^{\prime}}\bar{Q}(\bar{s}^{\prime},\bar{a}^{\prime})\big{]}$
@@
IV-D Joint state estimation
@@
where the conditional distribution $p(I_{1,t}\mid s_{2},s_{3})$ follows a multivariate Normal distribution with mean vector $\bm{\mu}_{1,t}$ and covariance matrix $\bm{\Sigma}_{1,t}$, and $\mathbf{b}_{1,t}$ denotes TX1’s belief vector at time $t$ – the probability that TX1 believes TX2 and TX3 are in states $s_{2}$ and $s_{3}$, respectively. Herein, the parameters $\bm{\mu}_{1,t}$ and $\bm{\Sigma}_{1,t}$ depend on the ARSS measurements up to time $t$.
@@
IV-E Local and joint Q-function updates
In Algorithm 1, the update of different $Q$-functions, as given in Table I depends on the transition direction between coordinated (C) and uncoordinated (U) states, as summarized in Table II. Individual $Q$-functions ($Q_{i}^{(n)}$ and $Q_{i}^{e}$) are updated in uncoordinated states, while the joint $Q$-function ($\bar{Q}$) is updated in coordinated states. In uncoordinated states, the joint $Q$-function is the sum of individual $Q$-functions since the TXs act independently.
In the coordinated states, TXs employ the following communication protocol: First, each TX independently estimates the joint state based on their local ARSS observations (as explained in Sec. IV-D) and shares the corresponding belief probability with the leader TX1. Next, TX1 selects the joint state with the highest belief probability and broadcasts the corresponding joint action. Each TX then executes its individual action and reports its costs and $Q$-functions to TX1, which subsequently updates the joint $Q$-function. This section of the algorithm is given in line 8.
@@ V Theoretical Analysis
We here provide several theoretical results for Algorithm 1.
@@ V-A Probabilistic Analysis
$Q_{t,i}^{(n)}(s_{i},a_{i})-Q^{*}_{i}(s_{i},a_{i})\sim D_{i,n}\left(0,\frac{\lambda_{i,n}^{2}}{3}\right),$ (10)
@@
This assumption is widely employed in RL literature: (i) it is used for non-MEMQ algorithms with $D_{i,n}$ being uniform or normal for $n=1$ case in [45, 17]; (ii) it is generalized to $n>1$ with no assumptions on $D_{i,n}$ for MEMQ algorithms in [7, 44, 5, 6]; and (iii) it is employed in MEMQ algorithms with $D_{i,n}$ being uniform in [37]. These assumptions are also theoretically and numerically justified in [7, 44, 5, 6]. This assumption is generalized to multiple agents in this work. We emphasize that this assumption is used solely to facilitate our theoretical analysis and may not hold exactly in practice. However, the assumption can be relaxed by allowing $D_{i,n}$ to have a non-zero mean while constraining a weighted sum of means to equal zero, which may be more feasible in practice [5]. Our theoretical results hold under both cases.
We herein leverage probabilistic results in [5], which were derived for the single-agent MEMQ to derive our results for multi-agent MEMQ. Hence, the proof techniques in [5] can be adapted here to prove these results. These results hold for any number of agents ($N_{T}$) and number of environments ($K_{i}$).
@@
$\mathbb{V}\Big{(}\bar{Q}_{t}(\bar{s},\bar{a})-\sum_{i=1}^{N_{T}}Q_{i}^{*}(s_{i},a_{i})\Big{)}\leq\sum_{i=1}^{N_{T}}\frac{f_{i}(\lambda_{i})}{K_{i}},$ (11)
where $\lambda_{i}=\max\limits_{n}\lambda_{i,n}$ is the estimation error variance of the worst environment for agent $i$, and $f_{i}$ is some function of the parameter $\lambda_{i}$.
This proposition has several implications. First, as the number of environments ($K_{i}$) increases, the upper bound on the estimation error variance decreases, which is consistent with the variance-reduction property of ensemble RL algorithms. Second, even if some agents employ a small number of environments, the estimation error variance can still be reduced by employing a larger number of environments for other agents, which differs from the single-agent case [5], where employing a smaller number of environments always lead to a larger estimation variance. Third, the function $f_{i}$ can be fine-tuned to further reduce the upper bound (see [5]). Finally, this result gives us the behavior of the estimation error variance in the finite $t$ case. Hence, the variance reduction property of Algorithm 1 holds even for small $t$ when the algorithm may be significantly far from convergence.
@@
$\lim_{t\rightarrow\infty}\mathbb{V}\Big{(}\bar{Q}_{t}(\bar{s},\bar{a})-\sum_{i=1}^{N_{T}}Q_{i}^{*}(s_{i},a_{i})\Big{)}\leq\frac{1-u}{1+u}\sum_{i=1}^{N_{T}}\lambda_{i}.$ (12)
@@
[…]

### v3

@@ III-2 State
The individual state of TXi is a triplet $s_{i}=(x_{i},y_{i},I_{i})$, where $(x_{i},y_{i})$ is its location with $x_{i},y_{i}\in[0,\Delta_{L},2\Delta_{L},...,L]$, and $I_{i}$ is the ARSS. Since ARSS may be continuous, we quantize $I_{i}$ to the nearest value in $I_{\text{min}},I_{\text{max}}]$, which is discretized into equal steps of size $\Delta_{I}$. The transition from $(x_{i},y_{i})$ to $(x_{i}^{\prime},y_{i}^{\prime})$ depends on the random movement of TXi, while the transition from $I_{i}$ to $I_{i}^{\prime}$ is influenced by several factors such as the random movement of all TXs, the distances between TXs, $L$, and $I_{thr}$. The joint state $\bar{s}$ is the concatenation of individual states as $\bar{s}=(s_{1},s_{2},...,s_{N_{T}})$, and is not available to the agents (i.e. agents do not have access to each others’ states).
The probability of transitioning from state $s_{i}$ to $s_{i}^{\prime}$ is:
$P(s_{i}^{\prime}\mid s_{i})=P_{x}(x_{i}^{\prime}\mid x_{i})\;P_{y}(y_{i}^{\prime}\mid y_{i})\;P_{I}(I_{i}^{\prime}\mid I_{i}),$ (5)
where $P_{x}(\cdot)$ and $P_{y}(\cdot)$ denote the probabilities of the user’s $x$- and $y$-coordinates performing a random walk over the possible directions. The term $P_{I}(\cdot)$ is the probability of transitioning from the ARSS value $I_{i}$ to $I_{i}^{\prime}$, which is influenced by the location of the users and their random movements, their local ARSS measurements, and the discretization and quantization of the continuous ARSS into the discrete representation.
The state-space is divided into two parts: coordinated and uncoordinated states. The joint state $\bar{s}$ is coordinated if there is at least one strong ARSS measurement in the wireless network, i.e., there is at least one agent $i$ such that $I_{i}>I_{thr}$, where $I_{thr}$ is the ARSS threshold; otherwise, it is uncoordinated. Thus, if at least one TX identifies the state as coordinated, it informs TX1, which then broadcasts this information to all TXs. Herein, high ARSS indicates an agent receives strong signals from nearby users, as signal strength decays with distance (path loss). Conversely, low ARSS occurs when the agent is isolated or far from other users, and receives weak signals. In real wireless networks, ARSS is analogous to perceived voice quality, which can degrade in dense networks due to interference and latency. In such high-density situations, corresponding to coordinated states, users can benefit from coordination via the central controller (i.e. the leader agent) by reducing interference, collisions, and latency.
We note that there are alternative approaches to determining coordinated states, such as based on predefined states [9] or proximity between agents [26]. Our approach is similar to [29], which also uses local RSS measurements to determine coordinated states. However, in our work, coordination is based on sufficiently high ARSS values. In addition, coordination is possible even if not all agents measure high ARSS levels. This makes our scheme robust against ARSS measurement inaccuracies in some agents, and scalable by reducing dependency on each agent’s measurement.
We denote $|\mathcal{S}_{C}|$ and $|\mathcal{S}_{U}|$ as the number of coordinated and uncoordinated states such that $|\mathcal{S_{C}}|+|\mathcal{S_{U}}|=|\mathcal{S}|$. We theoretically derive the optimal value of $I_{thr}$ that minimizes the lower bound on the probability of misdetection (i.e., detecting the state as coordinated while it is actually uncoordinated or vice versa) in the next section.
@@ III-3 Action
The action set of TXi is defined as
$\mathcal{A}_{i}=\{(m,b)\mid m\in\{0,1\},\;b\in\{1,2,\ldots,N_{B}\}\},$
where $m=0$ denotes staying stationary and $m=1$ denotes initiating movement. If $m=1$, the next position is sampled from a random walk. The index $b$ specifies the BS to which TXi connects; actions corresponding to BSs outside coverage are invalid and incur infinite cost. The joint action is $\bar{a}=(a_{1},a_{2},\ldots,a_{N_{T}})$ for each $a_{i}\in\mathcal{A}_{i}$. Herein, the action of each TXi directly affects its state transitions in (5).
@@ III-4 Cost function
$c_{i}=\beta_{1}\frac{P_{i}}{\log_{2}(1+\operatorname{SNR}_{i})}+\beta_{2}\mathcal{Q}(\sqrt{\text{SNR}_{i}})+\beta_{3}\mathcal{Q}(\frac{I_{thr}-I_{i}}{\sigma}),$ (6)
where $\operatorname{SNR}_{i}=\frac{P_{i}}{d_{i}^{2}(n+I_{i})}$, $d_{i}$ is distance between TXi and the BS it is connected to, $P_{i}$ is the transmission power of TXi which is fixed but may differ among agents, and $n$ is the zero-mean Gaussian noise as $n\sim\mathcal{N}(0,\sigma^{2})$. Herein, the first component is the efficiency cost. A higher SNR with constant transmitting power improves the effective data rate and reduces the efficiency cost. The second component is the reliability cost. The $\mathcal{Q}(x)=\frac{1}{2}\operatorname{erfc}\left(\frac{x}{\sqrt{2}}\right)$ represents the inverse relationship between SNR and bit error rate. Thus, higher SNR reduces the bit error rate and improves transmission reliability. The third component is the fairness cost. If the measured local ARSS exceeds the threshold $I_{thr}$, it suggests that some TXs are close to each other and may compete for the same resources (time slots, frequencies), potentially overloading some BSs while underutilizing others, which increases the fairness cost. The constants $\beta_{1},\beta_{2},\beta_{3}$ are the cost weights such that $\beta_{1}+\beta_{2}+\beta_{3}=1$. The joint cost is the sum of individual costs: $\bar{c}=\sum_{i=1}^{N_{T}}c_{i}(\bar{s}_{t},\bar{a}_{t})$. We emphasize that the individual cost function of each TX in (6) relies solely on local ARSS observations and does not depend on the knowledge of other TXs.
This cost function has several key properties. It is non-negative and bounded. It is highly non-linear and non-monotonic in some parameters (e.g., $P_{i}$), but monotonic in others (e.g., SNR and $I_{thr}$). Its non-convexity (due to the Gaussian $\mathcal{Q}$-function) presents optimization challenges. The effect of different costs and parameters can be controlled through the weights $\beta_{i}$.
@@ IV-A Preliminaries and initialization
Inputs: $l$, $\gamma$, $\alpha_{t}$, $\zeta_{t}$, $u_{t}$, $\mathbf{b}_{i,0}$, $Q_{i}^{(n)}$, $Q^{e}_{i}$, wireless parameters
@@
6:    Determine if current state is coordinated based on local ARSS, and if so, estimate it using the Bayesian approach
7:    Update the local and joint $Q$-functions as in Table II
8:  end for
@@
[…]
The inputs to Algorithm 1 are algorithm parameters (trajectory length $l$, discount factor $\gamma$, learning rate $\alpha_{t}$, exploration probability $\zeta_{t}$, update ratio $u_{t}$, initial belief vectors $\mathbf{b}_{i,0}$ for agent $i$ at $t=0$ (defined in Section IV-E), number of iterations $T$) and wireless network parameters ($L$, $\Delta_{L}$, $N_{T}$, $N_{B}$, $r_{j}$, $I_{min}$, $I_{max}$, $\Delta_{I}$, $\beta_{1},\beta_{2},\beta_{3}$, $P_{i}$, $k$, $I_{thr}$, $\sigma$). The outputs are the joint $Q$-table, $\bar{Q}$, and the corresponding joint policy, $\bar{\pi}$.
@@ IV-B Creating synthetic environments
Each TXi independently interacts with the real environment Env1, collects online samples $\{s_{i},a_{i},s_{i}^{\prime},c_{i}\}$ to update its Q-functions for the real environment, and stores this data in their buffer $\mathcal{B}_{i}$. Then, each TXi randomly samples offline mini-batches from $\mathcal{B}_{i}$ to update their individual estimated PTT for Env1 ($\mathbf{\hat{P}}_{i}$), and simultaneously builds and updates the PTT of its synthetic environment ($\mathbf{\hat{P}}_{i}^{n}$) by taking the matrix power of $\mathbf{\hat{P}}_{i}$ as explained in Sec.II-C. In this way, $Q$-functions are updated using online data, while offline samples from the buffer are used exclusively for kernel (PTT) estimation and updates. We emphasize that creating synthetic environments is not a one-time process; the PTTs of synthetic environments are continuously updated as each TX continues sampling from the real environment. Thus, the accuracy of the estimated synthetic environments improves over time.
@@
U $\rightarrow$ C $Q_{i}^{(n)}(s_{i},a_{i})\leftarrow(1-\alpha)Q_{i}^{(n)}(s_{i},a_{i})+\alpha\big{[}c_{i}(s_{i},a_{i})+\frac{\gamma}{3}\min_{\bar{a}^{\prime}}\bar{Q}(\bar{s}^{\prime},\bar{a}^{\prime})\big{]}$,  $Q^{e}_{i}(s_{i},a_{i})\leftarrow$ MEMQ
C $\rightarrow$ U $\bar{Q}(\bar{s},\bar{a})\leftarrow(1-\alpha)\bar{Q}(\bar{s},\bar{a})+\alpha\sum_{i=1}^{3}\big{[}c_{i}(\bar{s},\bar{a})+\gamma\min_{a_{i}^{\prime}}Q_{i}^{e}(s_{i}^{\prime},a_{i}^{\prime})\big{]}$
C $\rightarrow$ C $\bar{Q}(\bar{s},\bar{a})\leftarrow(1-\alpha)\bar{Q}(\bar{s},\bar{a})+\alpha\big{[}\sum_{i=1}^{3}c_{i}(\bar{s},\bar{a})+\gamma\min_{\bar{a}^{\prime}}\bar{Q}(\bar{s}^{\prime},\bar{a}^{\prime})\big{]}$
@@
IV-C Joint state estimation in Coordinated States
@@
where the conditional distribution $p(I_{1,t}\mid s_{2},s_{3})$ follows a multivariate Normal distribution with mean vector $\bm{\mu}_{1,t}$ and covariance matrix $\bm{\Sigma}_{1,t}$, and $\mathbf{b}_{1,t}$ denotes TX1’s belief vector at time $t$ – the probability that TX1 believes TX2 and TX3 are in states $s_{2}$ and $s_{3}$, respectively. Herein, the parameters $\bm{\mu}_{1,t}$ and $\bm{\Sigma}_{1,t}$ depend on the ARSS measurements up to time $t$. The computation of these parameters are given in the supplementary material.
@@
IV-D Local and Joint Q-function updates
In Algorithm 1, $Q$-function updates depend on transitions between coordinated (C) and uncoordinated (U) states. Hence, they are 4 different cases as shown in Table II.
In $\to$ U step, each agent independently updates its local $Q$-functions ($Q_{i}^{(n)},Q_{i}^{e}$). Agents update their local $Q$-functions across environments ($Q_{i}^{(n)}$) using standard $Q$-learning, and independently update their ensemble $Q$-functions ($Q_{i}^{e}$) via the single-agent MEMQ algorithm. To this end, several alternatives exist [6, 4, 5]. While these algorithms share similar goals, they differ in assumptions and network constraints. We adopt [4] as the single-agent MEMQ algorithm due to its broad applicability with minimal assumptions. For networks with specific constraints, one may instead use [6, 5]. We also assume that the joint $Q$-function satisfies an additivity property over the agents’ ensemble $Q$-functions as no information exchange takes place. Our formulation differs from prior work [39, 40, 41] in that we aggregate local ensemble $Q$-functions ($Q_{i}^{e}$) rather than traditional individual $Q$-functions ($Q_{i}^{(n)}$). To establish convergence of this update rule, one can directly apply the single-agent MEMQ convergence proof [4, 5] to each agent independently.
In $\to$ C step, the leader sums the agents reported local costs and uses this total to update $\bar{Q}$ from the previous iteration. In $\to$ U step, the leader agent transfers (backs up) each agent’s local ensemble $Q$-functions into $\bar{Q}$. As the system stays in coordination mode, individual agent-specific $Q$-functions are not updated. In $\to$ C step, the leader agent transfers $\bar{Q}$ into each agent’s $Q$-function, assigning an equal share of the expected discounted cost (e.g., $\tfrac{1}{3}$ for three agents). For these three steps, the system either operates centralized or with globally shared information. Thus, convergence follows standard $Q$-learning [31] by treating all agents as a single aggregated agent with global joint information.
When coordination is required, each TX first estimates the joint state from its local ARSS observations (Sec. IV-D) and shares the belief probability with the leader TX1. The leader then selects the joint state with the highest probability and broadcasts the joint action that minimizes the joint $Q$-function. Each TX executes its action and reports costs and $Q$-values to the leader, which then updates the joint $Q$-function. This section of the algorithm is given in line 7.
Similar update rules to Table II have been used in prior MARL studies [42, 39, 29, 9]. Our algorithm extends them by updating both the local individual $(Q_{i}^{(n)})$ and local ensemble $Q$-functions $(Q_{i}^{e})$.
@@ IV-D Local and Joint Q-function updates
We emphasize the partially decentralized nature of our algorithm and highlight key distinctions from fully centralized methods [27, 28, 43, 24]. In our approach, agents do not have access to the global joint state, whereas fully centralized methods assume complete global state observability. Second, the leader agent is active and computes the joint actions only when coordination is triggered (based on local ARSS measurements); otherwise, agents act independently without communication or information exchange. Communication is therefore sparse and conditional in contrast to continuous, full-network communication in centralized methods. As a result, our communication cost scales linearly with the number of agents, while centralized schemes often incur exponential growth with the joint state–action space.
@@ V Theoretical Analysis
We here provide several theoretical results for Algorithm 1. All proofs are given in the supplementary material.
@@ V-A Probabilistic Analysis
$\mathcal{X}^{(n)}_{t,i}(s_{i},a_{i})\!=\!Q_{t,i}^{(n)}(s_{i},a_{i})\!-\!Q^{*}_{i}(s_{i},a_{i})\!\sim\!D_{i,n}(0,\tfrac{\lambda_{i,n}^{2}}{3}),$ (11)
@@
This assumption is widely employed in RL literature: (i) it is used for non-MEMQ algorithms with $D_{i,n}$ being uniform or normal for $n=1$ case in [44, 11]; (ii) it is generalized to $n>1$ with no assumptions on $D_{i,n}$ for MEMQ algorithms in [38, 4, 5]; and (iii) it is employed in MEMQ algorithms with $D_{i,n}$ being uniform in [35]. These assumptions are also theoretically and numerically justified in [38, 4, 5]. This assumption is generalized to multiple agents in this work. We emphasize that this assumption is introduced solely to facilitate the theoretical analysis yet it can be relaxed as:
$\hskip-7.0pt\mathcal{X}^{(n)}_{t,i}(s_{i},a_{i})\!=\!Q_{t,i}^{(n)}(s_{i},a_{i})\!-\!Q^{*}_{i}(s_{i},a_{i})\!\sim\!D_{i,n}\left(\!\mu_{i,n},\!\tfrac{\lambda_{i,n}^{2}}{3}\!\right),\hskip-4.0pt$ (12)
with the condition that
$\sum_{n=1}^{K}w_{i}^{(n)}\mu_{i,n}=0,$ (13)
where $w_{i}^{(n)}$ denotes the weight of the $n^{\text{th}}$ synthetic environment in the MEMQ algorithm of agent $i$. This assumption is less restrictive than (11), as it only requires the weighted convex combination of biases to be zero, allowing individual means to be non-zero. Herein, the signs of $\mu_{i,n}$ depends on $n$. Thus, (i) over and underestimation effects cancel each other across different $n$, and (ii) higher-bias $Q$-functions receives smaller weights $w_{i}^{(n)}$ as shown in [4], ensuring the condition holds in practice. Our theoretical results hold under both cases. We also emphasize that both (11) and (12) assume that the $Q$-function estimation errors, not the $Q$-function updates themselves follow a random distribution with finite variance.
We herein leverage probabilistic results in [4], which were derived for the single-agent MEMQ to derive our results for multi-agent MEMQ. These results hold for any number of agents ($N_{T}$) and number of environments ($K_{i}$).
@@
$\mathbb{V}\Big{(}\bar{Q}_{t}(\bar{s},\bar{a})-\sum_{i=1}^{N_{T}}Q_{i}^{*}(s_{i},a_{i})\Big{)}\leq\sum_{i=1}^{N_{T}}\frac{f_{i}(\lambda_{i},t)}{K_{i}},$ (14)
where $\mathbb{V}$ is the variance operator, $\lambda_{i}=\max\limits_{n}\lambda_{i,n}$ is the estimation error variance of the worst environment for agent $i$, and $f_{i}$ is some function of the parameter $\lambda_{i}$.
This proposition has several implications. First, as the number of environments ($K_{i}$) increases, the upper bound on the estimation error variance decreases, which is consistent with the variance-reduction property of ensemble RL algorithms. Second, even if some agents employ a small number of environments, the estimation error variance can still be reduced by employing a larger number of environments for other agents, which differs from the single-agent case [4], where employing a smaller number of environments always lead to a larger estimation variance. Third, the function $f_{i}$ can be fine-tuned to further reduce the upper bound (see [4]). Finally, this result gives us the behavior of the estimation error variance in the finite $t$ case. We emphasize that this bound applies only to the variance, not the absolute error, between optimal and estimated $Q$-functions; hence, even if $K\to\infty$ drives the variance to zero, convergence to the optimal $Q$-function is not guaranteed in finite time $t$ due to the non-zero means in (12).
@@
$\lim_{t\rightarrow\infty}\mathbb{V}\Big{(}\bar{Q}_{t}(\bar{s},\bar{a})-\sum_{i=1}^{N_{T}}Q_{i}^{*}(s_{i},a_{i})\Big{)}\leq\frac{1-u}{1+u}\sum_{i=1}^{N_{T}}\lambda_{i}^{2}.$ (15)
@@
$\lim_{t\rightarrow\infty}\bar{Q}_{t}(\bar{s},\bar{a})\xrightarrow{\text{in mean square}}\sum_{i=1}^{N_{T}}Q_{i}^{*}(s_{i},a_{i}).$ (16)